Trace the curve x = a sin 2t (1+cos2t),
y=acos2t(1-cos2t), 0<=t<=π
1
Expert's answer
2020-05-08T17:24:24-0400

Let us determine the properties of the curve. First, we rewrite the formulae in form

"\begin{cases}
x(t) =2a\sin(2t)\cos^2t, \\
y(t) =2a\cos(2t)\sin^2t.
\end{cases}"

Or in form

"\begin{cases}
x(t) =4a\sin(t)\cos^3t, \\
y(t) =2a(1-2\sin^2t)\sin^2t.
\end{cases}"

We know that "\cos(\pi-t) = -\cos t , \;\; \sin(\pi-t) = \sin t". Therefore,

"\begin{cases}
x(\pi-t) = -x(t), \\
y(\pi-t) = y(t).
\end{cases}"

We see that the curve is is symmetrical about y-axis, and we should determine the shape of curve on "0≤t≤\dfrac{π}{2}" .


The curve passes through (0,0) if "t=0."


Let us determine maximum and minimum values of y. Minimum value of y will be at "t=\dfrac{\pi}{2}," because "\sin^2t" has the maximum value. So, "y(\frac{\pi}{2}) = -2a, \; x(\frac{\pi}{2})=0."


If we take a derivative of "y(t)" we get

"y'(t) = -4a\cos(t)\sin(t)(2\sin(t)-1)(2\sin(t)+1)".

On "0\le t \le \dfrac{\pi}{2}" extrema are located at "0, \dfrac{\pi}{2}" (see above) and "\dfrac{\pi}{6}" (maximum, "x=\dfrac{3\sqrt3}{4}a, \; y = \dfrac{a}{4}" ), because at these points "y'(t)=0."


Next let us determine maximum and minimum values of x on "0≤t≤\dfrac{π}{2}" . We know that the curve passes through (0,0), so minimum value of x is 0. If we take the derivative of "x(t)", we get

"x'(t) = 4a\cos^2t(\cos^2t - 3\sin^2t) = 4a\cos^2t(1-4\sin^2t)."

"x'(t)=0" at "t=\dfrac{\pi}{2}" (minimum, x=0) and "t=\dfrac{\pi}{6}." We see that for "t=\dfrac{\pi}{6}" both x and y have their maximum values.


Next, let us determine "x" values when "y=0." We can see that "y(t)=0" if "\sin t = 0" or "\sin^2 = \dfrac12." On "0\le t \le \dfrac{\pi}{2}" this means that "t=0" or "t=\dfrac{\pi}{4}." If "t=0" , we get "x=y=0" . If "t=\dfrac{\pi}{4}" , we get "x=a."


Now let us write points we have calculated above:

"x \quad \quad \quad y \\
-----\\
0 \quad \quad \quad 0 \\
 a \quad \quad \quad 0\\
\dfrac{3\sqrt3}{4}a \quad \dfrac{a}{4} \\
0 \quad \quad -2a"


We may also calculate the derivative "\dfrac{dy}{dx} = \dfrac{-4a\cos t\sin t(4\sin^2t-1)}{4a\cos^2t(1-4\sin^2t)} = \tan t." The derivative if positive on "0\le t \le \dfrac{\pi}{2}" and it is an increasing function and equals to 0 at "(x,y) = (0,0)." We'll use this information when creating a graph.